# Enhancing Domain-Specific Encoder Models with LLM-Generated Data: How to Leverage Ontologies, and How to Do Without Them

### v1

@@ 4.3 Evaluation Datasets
The Hypothesis Classification task Brinner et al. (2022) is a 10-class classification task on identifying which of 10 major hypotheses in invasion biology is addressed in a given scientific abstract. Due to class imbalance, we report both micro F1 and macro F1 scores.
The Hypothesis Span Prediction task Brinner et al. (2024) is a token-level prediction task based on the same abstracts as the INAS classification task. Annotators provide span-level evidence annotations for each hypothesis and we evaluate the model’s ability to predict the tokens that were annotated (Token F1) as well as the ability to recognize complete spans (Span F1).
The EICAT Impact Classification task Brinner and Zarrieß (2025) is a classification task on assessing the impact of an invasive species as reported in a given scientific full text, assigning it to one of six predefined impact categories. We evaluate performance using macro F1 and micro F1 scores.
The EICAT Impact Evidence task Brinner and Zarrieß (2025) leverages evidence annotations provided by the EICAT classification dataset, created by domain experts who identified sentences in the full-texts indicating the species’ impact category. We evaluate the model’s ability to rank relevant sentences within a full text using the normalized discounted cumulative gain (NDCG) metric.
@@
Since we observed a high variance between results for different training runs, we train 7 models for the hypothesis and impact classification tasks and 3 models for the remaining tasks, thus reporting average performance on the test sets. For details on task setup, dataset sizes and training methodologies, please refer to Appendix A.
@@ 4.4 Results
First, we observe that traditional MLM pretraining on sentences extracted from just 5,000 scientific abstracts yields substantial performance improvements across all tasks compared to the standard DeBERTa model, raising the benchmark score from 0.483 to 0.507.
As a baseline, we also assess the impact of MLM pretraining on synthetic definitions. While this also resulted in increased performance, the gains are smaller than those achieved through pretraining on abstract sentences. Additionally, despite the datasets being of similar size, optimal performance with ontology definitions is reached after approximately 40,000 batches, in contrast to 200,000 batches for MLM on abstract sentences, which is analyzed further in Section 6.2.
As a last MLM baseline, we investigate MLM pretraining on a mixture of synthetic definitions and abstract sentences. Since initial experiments using a 1:1 ratio led to worse results compared to training on abstract sentences alone, we adjusted the ratio to 1:3 (ontology definitions to abstract sentences), resulting in improved performance compared to using abstract sentences alone, suggesting that concept definitions provide useful additional information to the model.
@@
The most notable improvements arise when combining SIM pretraining on synthetic ontology definitions with MLM pretraining on abstract sentences. This approach leads to substantial performance gains across most tasks compared to MLM pretraining alone. Specifically, the overall benchmark score increases from 0.507 (MLM on abstract sentences) to 0.538. Notably, the substantial improvement over using either pretraining method individually suggests a synergistic effect, indicating that SIM pretraining enhances the understanding of individual concepts, while MLM pretraining strengthens the model’s grasp of relationships between concepts and general language understanding. As a result, this combined approach outperforms models trained on millions of abstracts from the broader biomedical domain, such as PubMedBERT Jeong and Kim (2022) and SciDeBERTa Kim et al. (2023), which generally are strong baselines in this field Brinner et al. (2022).
@@ 5 Using LLM-Extracted Keywords
In the previous section, we explored the performance improvements achieved by combining our proposed contrastive loss on ontology-derived data with traditional MLM pretraining. While this approach is highly valuable in domains with available ontologies, many fields may lack such structured resources. To address this limitation, we explore the feasibility of using an LLM for constructing a dataset of domain-relevant concepts, definitions, and relations using only a small set of scientific abstracts. We compare results achieved on our original dataset of 5,000 abstracts with those using ontology-derived data and also evaluate how well our approach scales with increasing dataset size.
@@ 5.1 Dataset Creation
    Keyword Extraction: We extract domain-relevant concepts in the form of keywords from scientific abstracts using LLaMA-3-8B Grattafiori et al. (2024). This is achieved by appending the string "Keywords:" to each abstract and allowing the language model to generate a continuation, effectively identifying key concepts within the text.
@@ 5.2 Results
We first evaluate traditional MLM pretraining on keyword definitions derived from LLM-extracted keywords, which leads to slight performance improvements over the standard DeBERTa base model (score: 0.483) by achieving scores of 0.492 when trained solely on keyword definitions and 0.507 when combined with abstract sentences. However, these gains are less pronounced than those using LLM-generated definitions for ontological concepts, indicating that ontological concepts offer more valuable information to the encoder model (compare Section 6).
In contrast, SIM pretraining on keyword definitions yields slightly better performance than using ontology definitions. This advantage may be attributed to dataset size, as the LLM extracted 23,597 unique keywords from the abstracts, compared to 5,179 concepts from the ontologies. Notably, this enhanced performance lets SIM pretraining on data extracted from 5,000 abstracts surpass MLM pretraining on the same abstracts. This finding not only validates our proposed pretraining approach but also suggests that the LLM has enriched our base dataset with valuable information.
However, we observe a reverse trend when examining the combination of MLM pre-training on abstract sentences and SIM pretraining on synthetic definitions. Here, leveraging ontology data results in a significantly greater performance boost than using keywords definitions, which we analyze further in Section 6. Still, the resulting model using just 5,000 abstracts outperforms SciDeBERTa, which was trained on millions of scientific abstracts.
Lastly, we assess the effect of varying dataset sizes on our pretraining pipeline. While an increase in data availability leads to more detected keywords for SIM pretraining, it also leads to more abstract sentences for MLM pretraining. This may diminish the relative value added by the LLM. However, as shown in Table 2, even with larger datasets, our fully automated knowledge injection strategy consistently outperforms traditional MLM pretraining, even though both are based on the same dataset.
Despite efforts to mitigate variance by training multiple models per task, we also note that individual results remain subject to fluctuation. Therefore, we consider the average scores across all dataset sizes - 0.517 for MLM pretraining and 0.529 for combined pretraining - as the most reliable indicators of the substantial performance improvements achievable with our pipeline.
@@ 6 Discussion
6.1 Are Ontologies Replacable?
Our experiments demonstrate that injecting domain-specific knowledge from ontologies into encoder models can substantially enhance downstream performance. Interestingly, we also found that ontological knowledge can - to some extent - be replaced by a combination of automatically extracted keywords, definitions, and co-occurrence statistics. While this might suggest that ontologies add little value beyond these extracted elements, we argue that this conclusion is premature:
@@ 6.1 Are Ontologies Replacable?
Also, we point out that a combination of synthetic data and abstract sentences leads to superior results when ontology-based definitions are used instead of keyword definitions (both for MLM and SIM). This disparity may stem from the fact that information extracted from the abstracts is inherently tied to the same dataset, thus offering less additional insight compared to the disconnected and therefore more informative ontology.
@@ 6.2 Investigating Model Collapse
Previous studies have identified a risk of model collapse when training on generated data (see Section 2). Similarly, in our experiments, we observed that both MLM and SIM training on synthetic data reached peak performance after approximately 40,000 batches, after which performance began to decline. In contrast, training on the dataset consisting of abstract sentences peaked at around 200,000 batches, with performance remaining stable even when training for twice as long. This suggests that while the generated data provides valuable information, excessive use can still lead to model collapse.
@@
Interestingly, we found that performance degradation was much less pronounced for SIM training than for MLM training on synthetic definitions. This is likely due to much weaker gradient signals after the peak has been reached, as most training triples eventually reach zero loss. This has the positive effect that, when SIM pretraining on synthetic data is combined with MLM training on abstract sentences, the risk of model collapse is effectively mitigated because the gradients from SIM training are not strong enough to induce this effect.
This is in contrast to MLM training on a combination of abstract sentences and synthetic definitions, for which performance declined compared to training on abstract sentences alone when both sources of data were used in equal proportion. This suggests that in this setting, the signal leading to model collapse is too strong, leading us to adopt a 1:3 ration in our experiments. These findings highlight the advantage of our proposed pretraining scheme over traditional MLM, as it enables effective utilization of synthetic data while avoiding detrimental effects on model stability.
@@ 7 Conclusion
In this study, we investigated the use of LLM-generated, synthetic data for continual pretraining of domain-specific encoder models. The approach demonstrates how to utilize domain specific ontologies or derive domain information through LLM-extraction from scientific abstracts for domains where ontologies may not be available.
@@ 8 Limitations
We note several limitations of our approach: First, while we demonstrate strong performance in the domain of invasion biology, its applicability to other domains remains uncertain and requires further evaluation.
@@ A.1 Data Generation
Concept relations are identified using co-occurrence counts as described in Section 5.1. For the dataset consisting of 5000 abstracts, we treat concepts as related if they co-occur in at least 5 abstracts, with this number being increased by one for each increase in dataset size. Since many concepts do not occur that often, this lead to each concept being related to about 0.5 other concepts.
@@ A.2 Model Training
We evaluate various pretraining strategies. Initially, we selected the optimal model checkpoint based on validation loss; however, we found that training for significantly longer improved downstream performance, even when the validation loss did not decrease. For this reason, we adopted a strategy of saving model checkpoints at different epochs and evaluating them on the INAS classification task, thus identifying the number of batches that are optimal for a given pretraining method. Once established, we retrained the final models used in our evaluation from scratch using the predetermined number of epochs.
@@ A.3.1 INAS Classification
The INAS classification task Brinner et al. (2022) is a 10-class classification problem, where the goal is to determine which of 10 prominent hypotheses are addressed in a given scientific abstract. We use the updated labels provided by Brinner et al. (2024). The task is a multi-label classification task, meaning that multiple hypotheses can be addressed within a single abstract.
@@ A.3.2 INAS Span Prediction
The INAS Span Prediction task Brinner et al. (2024) is closely related to the INAS classification task and is based on the same dataset. However, instead of classifying abstracts, it involves identifying spans of text indicative of the 10 hypotheses, as annotated by human experts.
@@
For further details on these metrics, see Brinner et al. (2024).
@@ A.3.3 EICAT Classification
The EICAT classification task Brinner and Zarrieß (2025) is concerned with classifying the ecological impact of an invasive species as reported in a scientific full-text paper. The categories include five different impact levels plus a “Data Deficient” category, resulting in a six-class classification problem.
@@
Since most encoder models cannot process entire full-texts at once, Brinner and Zarrieß (2025) explored strategies for selecting relevant sentence subsets for training and evaluation. One effective and unbiased approach is the selection of random sentences, which we adopt. During testing, each model receives 20 different random sentence selections per paper, with the final classification determined via majority voting.
@@ A.3.4 EICAT Evidence Selection
The EICAT evidence selection task Brinner and Zarrieß (2025) is a binary sentence classification problem. While annotating scientific full-texts for the EICAT classification task, human experts identified key sentences that served as evidence for impact assessments. The goal of this task is to predict whether a given sentence is evidence for an EICAT impact assessment.
@@ A.4 Evaluation Details
Due to the high variance in the model’s predictions, we train 7 models for the INAS classification and EICAT classification tasks, as well as 3 models for the other tasks that take significantly longer for each training run. Final results are reported as the average performance across all runs. To compute a final benchmark score, we first average the performance metrics for each task separately and then compute an overall average across all tasks.
For both EICAT-related tasks, we observed occasional training runs (across all pretraining types) where models exhibited drastically lower performance, often predicting only a single class for all samples. We attribute this to the dataset’s extreme class imbalance, that, for some random seeds, leads to degenerate states that the model is unable to escape. In such cases, training runs were repeated to avoid reporting results that reflect random failures rather than actual model performance.

### v2

@@ 4.3 Evaluation Datasets
The Hypothesis Classification task brinner2022linking is a 10-class classification task on identifying which of 10 major hypotheses in invasion biology is addressed in a given scientific abstract. Due to class imbalance, we report both micro F1 and macro F1 scores.
The Hypothesis Span Prediction task brinn2024weaklyclaim is a token-level prediction task based on the same abstracts as the INAS classification task. Annotators provide span-level evidence annotations for each hypothesis and we evaluate the model’s ability to predict the tokens that were annotated (Token F1) as well as the ability to recognize complete spans (Span F1).
The EICAT Impact Classification task brinner2025efficient is a classification task on assessing the impact of an invasive species as reported in a given scientific full text, assigning it to one of six predefined impact categories. We evaluate performance using macro F1 and micro F1 scores.
The EICAT Impact Evidence task brinner2025efficient leverages evidence annotations provided by the EICAT classification dataset, created by domain experts who identified sentences in the full-texts indicating the species’ impact category. We evaluate the model’s ability to rank relevant sentences within a full text using the normalized discounted cumulative gain (NDCG) metric.
@@
To mitigate variance inherent to model training, we train 7 models for the hypothesis and impact classification tasks and 3 models for the remaining tasks and report the average performance. For details on task setup, dataset sizes and training methodologies, please refer to Appendix A.
@@ 4.4 Results
First, we observe that traditional MLM pretraining on sentences extracted from just 5,000 scientific abstracts yields significant improvements across all tasks compared to the DeBERTa baseline, raising the benchmark score from 0.483 to 0.507.
As a baseline, we also assess the impact of MLM pretraining on synthetic definitions. While this also resulted in increased performance, the gains are smaller than those achieved through pretraining on abstract sentences. Additionally, despite the datasets being of similar size, optimal performance with synthetic definitions is reached after approximately 40K batches, in contrast to 200K batches for MLM on abstract sentences, which is analyzed further in Section 6.2.
As a last MLM baseline, we investigate MLM pretraining on a mixture of synthetic definitions and abstract sentences. Since initial experiments using a 1:1 ratio led to worse results compared to training on abstract sentences alone, we adjusted the ratio to 1:3 (ontology definitions to abstract sentences), resulting in improved performance compared to using abstract sentences alone and suggesting that concept definitions provide useful additional information to the model.
@@
The most notable improvements arise when combining SIM pretraining on synthetic ontology definitions with MLM pretraining on abstract sentences. This approach leads to substantial performance gains across most tasks compared to MLM pretraining alone. Specifically, the overall benchmark score increases from 0.507 (MLM on abstract sentences) to 0.538. Notably, the substantial improvement over using either pretraining method individually (or over using the combined data for MLM) suggests a synergistic effect, indicating that SIM pretraining enhances the understanding of individual concepts, while MLM pretraining strengthens the model’s grasp of relationships between concepts and general language understanding. As a result, this combined approach outperforms models trained on millions of abstracts from the broader biomedical domain, such as PubMedBERT Gu_2021 and SciDeBERTa 9791256, which generally are strong baselines in this field brinner2022linking.
@@ 5 Using LLM-Extracted Keywords
In the previous section, we explored the performance improvements achieved by combining our proposed similarity loss on ontology-derived data with traditional MLM pretraining. While this approach is highly valuable in domains with available ontologies, many fields may lack such structured resources. To address this limitation, we explore the feasibility of using an LLM for constructing a dataset of domain-relevant concepts, definitions, and relations using only a small set of scientific abstracts. We compare results achieved on our original dataset of 5,000 abstracts with those using ontology-derived data and also evaluate how well our approach scales with increasing dataset size.
@@ 5.1 Dataset Creation
    Keyword Extraction: We extract domain-relevant concepts in the form of keywords from scientific abstracts using LLaMA-3-8B grattafiori2024llama3herdmodels. This is achieved by appending the string "Keywords:" to each abstract and allowing the language model to generate a continuation, effectively identifying key concepts within the text.
@@ 5.2 Results
We again evaluate MLM pretraining on the new dataset of LLM-generated keyword definitions as a baseline, which leads to slight improvements over the standard DeBERTa base model by achieving scores of 0.492 when trained solely on keyword definitions and 0.507 when combined with abstract sentences. However, these gains are less pronounced than those using LLM-generated definitions for ontological concepts, indicating that ontological concepts offer more valuable information to the encoder model (compare Section 6).
In contrast, SIM pretraining on keyword definitions yields slightly better performance than using ontology definitions, which may be attributed to dataset size as the LLM extracted 23,597 unique keywords from the abstracts, compared to 5,179 concepts from the ontologies. Notably, this lets SIM pretraining on data extracted from 5,000 abstracts outperform MLM pretraining on that same dataset, thus validating our proposed pretraining approach and suggesting that the LLM has enriched our base dataset with valuable information.
Combining SIM and MLM pretraining again leads to improved results compared to either strategy alone, thus undermining the synergistic effects. However, the performance gains are weaker than those achieved using the ontology-derived data (0.520 vs. 0.538), which we analyze further in Section 6. Still, the resulting model using just 5,000 abstracts outperforms SciDeBERTa, which was trained on millions of scientific abstracts.
Lastly, we assess the effect of varying dataset sizes on our pretraining pipeline. While an increase in data availability leads to more detected keywords for SIM pretraining, it also leads to more abstract sentences for MLM pretraining, which may diminish the relative value added by the LLM. However, as shown in Table 2, even with larger datasets, our fully automated knowledge injection strategy consistently outperforms traditional MLM pretraining, even though both are based on the same dataset.
Despite efforts to mitigate variance by training multiple models per task, individual results still remain subject to fluctuation (see Appendix A.6 for an analysis on statistical significance). Therefore, we consider the average scores across all dataset sizes - 0.517 for MLM pretraining and 0.529 for combined pretraining - as the most reliable indicators of the substantial performance improvements achievable with our pipeline.
@@ 6 Discussion
6.1 Are Ontologies Replaceable?
Our experiments demonstrate that injecting domain-specific knowledge from ontologies into encoder models can substantially enhance downstream performance. Notably, we also found that this knowledge can - to some extent - be replaced by a combination of LLM-extracted keywords, definitions, and co-occurrence statistics. Still, we argue that ontologies are a more valuable resource, which is supported by several observations.
@@ 6.1 Are Ontologies Replaceable?
Second, we find that a combination of synthetic data and abstract sentences leads to superior results when ontology-based definitions are used instead of keyword definitions (both for MLM and SIM). This disparity may stem from the fact that information extracted from the abstracts is inherently tied to the same dataset, thus offering less additional insight compared to the disconnected and therefore more informative ontology.
@@ 6.2 Investigating Model Collapse
Previous studies have identified a risk of model collapse when training on LLM-generated data (see Section 2). Similarly, we observed that both MLM and SIM training on synthetic data reached peak performance after approximately 40K batches, after which performance began to decline. In contrast, training on the dataset consisting of abstract sentences peaked at around 200K batches, with performance remaining stable even when training for twice as long. This suggests that while the generated data provides valuable information, excessive use can still lead to model collapse.
@@
Nevertheless, we found that performance degradation with synthetic data was much less pronounced for SIM training compared to MLM. This is likely due to weaker gradient signals after the peak has been reached, as most training triples eventually reach zero loss. This has the positive effect that, when SIM pretraining on synthetic data is combined with MLM training on abstract sentences, the risk of model collapse is effectively mitigated because the weak (but still informative) gradients from SIM training are not strong enough to induce this effect.
This is in contrast to MLM training on a combination of abstract sentences and synthetic definitions. Here, performance declined compared to training on abstract sentences alone when both sources of data were used in equal proportion. This suggests that in this setting, the signal leading to model collapse is too strong, leading us to adopt a 1:3 ration in our experiments.
Ultimately, these findings highlight the advantage of our proposed pretraining scheme over traditional MLM, as it enables effective utilization of synthetic data while avoiding detrimental effects on model stability.
@@ 7 Conclusion
In this study, we investigated the use of LLM-generated, synthetic data for continual pretraining of domain-specific encoder models, demonstrating how to utilize domain specific ontologies or derive domain information through LLM-extraction from scientific abstracts for domains where ontologies may not be available.
@@ 8 Limitations
We note several limitations of our approach: First, while we demonstrate strong performance in the domain of invasion biology, its applicability to other domains remains uncertain and requires further evaluation, which was not possible to include within this study given the extend of the existing evaluations and analyses.
@@
9 Acknowledgments
This work was funded by the European Regional Development Fund within the project "LLM4KMU - Optimierter Einsatz von Open Source Large Language Modellen in KMU", and was partially produced within the focus group "Mapping Evidence to Theory in Ecology: Addressing the Challenges of Generalization and Causality" at the Center for Interdisciplinary Research, Bielefeld.
@@ Appendix A Experimental Details
Code for training and evaluation, training datasets and the best-performing encoder model checkpoint are available at github.com/inas-argumentation/Ontology_Pretraining.
@@ A.1 Data Generation
Concept relations are identified using co-occurrence counts as described in Section 5.1. For the dataset consisting of 5000 abstracts, we treat concepts as related if they co-occur in at least 5 abstracts, which we selected manually by observing and assessing exemplary related concepts. Since many concepts occur rarely, this lead to each concept being on average related to about 0.5 other concepts. For larger dataset sizes, we adjust the number of co-occurrences that are required for two concepts to be related so that the number of related concepts for each concept stays roughly constant at 0.5, thus leading to a comparable assessment.
@@ A.2 Model Training
We evaluate various pretraining strategies. Initially, we selected the optimal model checkpoint based on validation loss; however, we found that training for significantly longer improved downstream performance, even when the validation loss did not decrease. For this reason, we adopted a strategy of saving model checkpoints at different epochs and evaluating them on the INAS classification task. We then used this evaluation to identify the number of batches that are optimal for a given pretraining method. Once this number is established, we retrained the final models used in our evaluation from scratch using the predetermined number of epochs.
@@ A.3.1 INAS Classification
The INAS classification task brinner2022linking is a 10-class classification problem, where the goal is to determine which of 10 prominent hypotheses are addressed in a given scientific abstract. We use the updated labels provided by brinn2024weaklyclaim. The task is a multi-label classification task, meaning that multiple hypotheses can be addressed within a single abstract.
@@ A.3.2 INAS Span Prediction
The INAS Span Prediction task brinn2024weaklyclaim is closely related to the INAS classification task and is based on the same dataset. However, instead of classifying abstracts, it involves identifying spans of text indicative of the 10 hypotheses, as annotated by human experts.
@@
For further details on these metrics, see brinn2024weaklyclaim.
@@ A.3.3 EICAT Classification
The EICAT classification task brinner2025efficient is concerned with classifying the ecological impact of an invasive species as reported in a scientific full-text paper. The categories include five different impact levels plus a “Data Deficient” category, resulting in a six-class classification problem.
@@
Since most encoder models cannot process entire full-texts at once, brinner2025efficient explored strategies for selecting relevant sentence subsets for training and evaluation. One effective and unbiased approach is the selection of random sentences, which we adopt. During testing, each model receives 20 different random sentence selections per paper, with the final classification determined via majority voting.
@@ A.3.4 EICAT Evidence Selection
The EICAT evidence selection task brinner2025efficient is a binary sentence classification problem. While annotating scientific full-texts for the EICAT classification task, human experts identified key sentences that served as evidence for impact assessments. The goal of this task is to predict whether a given sentence is evidence for an EICAT impact assessment.
@@ A.4 Evaluation Details
Due to the variance inherent to training models on evaluation tasks, we train 7 models for the INAS classification and EICAT classification tasks, as well as 3 models for the other tasks that take significantly longer for each training run. Final results are reported as the average performance across all runs. To compute a final benchmark score, we first average the performance metrics for each task separately and then compute an overall average across all tasks.
For some tasks, we observed occasional training runs (across all pretraining types) where models exhibited drastically lower performance caused by degenerate states that only predict a single class for all samples. We attribute this to the dataset’s extreme class imbalance, that, for some random seeds, leads to degenerate states that the model is unable to escape. In such cases, training runs were repeated to avoid reporting results that reflect random failures rather than actual model performance.
@@ A.5 Ablation
A.6 Statistical Significance
We perform multiple runs for each task to reduce variance in our reported results. To fully undermine our key results, we perform a permutation-based statistical significance test that takes all 20 (or 80 for the multiple dataset sizes) individual results that contribute to the final benchmark score into account. According to this, the following results are statistically significant ($p<0.05$):
  • The superiority of all pretraining methods (except MLM pretraining using just keyword definitions) over DeBERTa base.
  • The superiority of combining MLM pretraining on abstract sentences with similarity pretraining on ontology definitions compared to just MLM pretraining on abstract sentences.
  • The superiority of combining MLM pretraining on abstract sentences with similarity pretraining on keyword definitions if evaluated over all dataset sizes.
  • The superiority of MLM+SIM pretraining using ontology data over MLM+SIM pretraining using abstract-derived keyword data.
Thus, the following key insights are supported by statistical significance:
  • SIM pretraining alone is a valid pretraining strategy that improves performance.
  • Combining SIM pretraining with MLM pretraining leads to improved results compared to just MLM pretraining alone. This holds both for the ontology-based and LLM-extracted keyword-based data.
  • Ontology data is a more valuable resource than data reliant on LLM extracted keywords.